### v3

# Cosmological phase-space analysis of $f(G)$-theories of gravity

Abstract
The impact of topological terms that modify the Hilbert-Einstein action is here explored by virtue of a further $f(G)$ contribution. In particular, we investigate the phase-space stability and critical points of an equivalent scalar field representation that makes use of a massive field, whose potential is function of the topological correction. To do so, we introduce to the gravitational Action Integral a Lagrange multiplier and model the modified Friedmann equations by virtue of new non-dimensional variables. We single out dimensionless variables that permit a priori the Hubble rate change of sign, enabling regions in which the Hubble parameter either vanishes or becomes negative. In this respect, we thus analyze the various possibilities associated with a universe characterized by such topological contributions and find the attractors, saddle points and regions of stability, in general. The overall analysis is carried out considering the exponential potential first and then shifting to more complicated cases, where the underlying variables do not simplify. We compute the eigenvalues of our coupled differential equations and, accordingly, the stability of the system, in both a spatially-flat and non-flat universe. Quite remarkably, regardless of the spatial curvature, we show that a stable de Sitter-like phase that can model current time appears only a small fraction of the entire phase-space, suggesting that the model under exam is unlikely in describing the whole universe dynamics, i.e., the topological terms appear disfavored in framing the entire evolution of the universe.
Gauss-Bonnet; $f(G)$-gravity; Equilibrium points
I Introduction
Challenging General Relativity has acquired much more attention in view of the most recent cosmological observations that certify the existence of a dark sector 2006IJMPD..15.1753C ; Gruber:2013wua ; delaCruz-Dombriz:2016bqh ; Sahni:2006pa ; Dunsby:2016lkw ; Luongo:2025iqq , fully-unknown, albeit necessary for both clustering and accelerating the universe Bamba:2012cp ; 2001IJMPD..10..213C ; 2003PhRvL..90i1301L . Accordingly, conceptual limitations or experimental drawbacks are becoming gradually more explored Simon2005 , especially in strong gravitational fields 2020A&A…641A.174L ; 2021MNRAS.503.4581L , where gravity breaks down Kiefer:2023bld or a non-trivial mechanism of vacuum energy production is expected Martin:2012bt leading to the cosmological constant problem Luongo:2018lgy ; Belfiglio:2022qai ; Luongo:2023aaq and so on.
The advantage of extending the Hilbert-Einstein’s action lies on the fact that Einstein’s theory is not renormalizable 2003PhR…380..235P ; 1992ARA&A..30..499C and, in turn, cannot be conventionally quantized. The cosmological constant problem, for example, remains an open challenge of our current understanding weinberg1989cosmological ; Luongo:2018lgy ; Capozziello:2025bsm and limits the possibility to characterize the cosmological constant today, as experimentally found 2024A&A…690A..40L ; Khadka:2021vqa ; Luongo:2021nqh ; Dunsby:2015ers ; Vilardi:2024cwq ; Bamba:2012cp ; Capozziello:2021xjw .
To face the problem of renormalization at one loop requires the Einstein-Hilbert action to be supplemented with additional higher-order curvature terms that, mostly turn out to be not unitary. Analogously, when quantum corrections, for example under the form of string modifications, are included, the effective low-energy gravitational action naturally incorporates higher-order curvature invariants, leading to experimental signature in extremely strong gravity regimes such the cosmological Planck scale, black hole singularities, etc.
However, signatures of this picture may be found even at the level of infrared scales, providing evidence in favor of modifying Einstein’s gravity at all scales.
A natural extension of Einstein’s gravity may lead to $f(R)$ theories of gravity Sotiriou:2008rp ; Nojiri:2008nk ; Nojiri:2008nt , constructed as analytic functions of the Ricci scalar $R$, whose prototype is offered by a second order term, $\propto R^{2}$, named scalaron Starobinsky:1980te . Viable $f(R)$ models, fulfilling both cosmological and local gravity constraints, have not really been found so far, and, in general, the matter sector is mainly coupled with such theories, limiting the consistency with observations.
Going through these paradigms, in view of the large number of proposals, one can find theories with a Lagrangian density being a combination of terms Barrow:1988xh ; Barrow:1991hg ; Bueno:2016dol $R_{\mu\nu}R^{\mu\nu}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$, namely superpositions of the Ricci and Riemann tensors, respectively, that are jeopardized by unphysical spin-2 ghost instabilities. We mention here, for the sake of completeness, other approaches such as torsion Nesseris:2013jea ; Geng:2011aj ; Paliathanasis:2016vsw ; Duchaniya:2023aeu ; DAgostino:2018ngy or non-metricity Heisenberg:2023lru ; Koussour:2023rly ; Solanki:2022ccf ; Anagnostopoulos:2021ydo that can, as well, provide hints on how to facethe aforementioned issue, see e.g. BeltranJimenez:2019esp .
In this respect we single out the study of Gauss-Bonnet (GB) terms Padmanabhan:2013xyr ; Dotti:2007az ; Charmousis:2002rc ; Pozdeeva:2019agu , defined as, $G\equiv R^{2}-4R^{\mu\nu}R_{\mu\nu}+R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}$, that can significantly improve the above picture. The GB scalar is the first topological scalar provided within the Lovelock theory of gravity Lovelock:1971yv ; Lovelock:1972vz . The latter theory is the second-order alternative to General Relativity gravitational theory in higher-dimensional geometries. However, in the case of a four-dimensional manifold Lovelock’s topological scalars are reduced topopological boundary terms, and general relativity is recovered. Thus, in order to understand the effects of these scalars in low-dimensional geometries, nonlinear terms are introduced in the gravitational action integral.
Over the past decades, various studies have explored this topic, generally revealing that satisfying local gravity constraints DeFelice:2009aj ; Yousaf:2019upm ; Bhatti:2020det remains challenging when the GB term is responsible for dark energy. However, this appears feasible for certain plausible $f(G)$ Lagrangians Nojiri:2005jg that yield a cosmological constant Li:2007jm ; Nojiri:2021mxf ; Nojiri:2024nlx ; Myrzakulov:2024sne ; Lohakare:2024ize ; Bamba:2017cjr ; Nojiri:2010wj ; Nojiri:2017ncd ; Nojiri:2018ouv ; Vernov:2021hxo .
Although a definitive conclusion on the validity of extended gravity, and specifically $f(G)$ models, has yet to be reached, investigating their stability and properties remains crucial for assessing their potential role in cosmology.
Motivated by the above considerations, we here focus on $f(G)$ theories of gravity and analyze the corresponding stability and phase-space dynamics. To do so, we limit our study to a FLRW background, reconsidering $f(G)$ theories as scalar field-equivalent representation, modifying the corresponding Lagrangian by including a Lagrangian multiplier. After computing the modified Friedmann equations, we formulate the cosmological dynamics by virtue of appropriate dynamical variables, rewritten in order to account for all possible behaviors of the universe, namely enabling the Hubble parameter to change sign, as a consequence of the dynamical behaviour. The corresponding autonomous system of dynamical equations has been obtained and the computation of eigenvalues has been found. Analyzing then the phase-space we obtain the existence of attractors and explore the stability of each point under examination. Following that the stability at asymptotic regime also has been found. To do so, we employ a particular version of power-law potential and compare our findings with previous works. As a matter of fact, we emphasize that the so-constructed model converges to a de Sitter phase
The paper is organized as follows. In Sect. II, we introduce the concepts related to $f(G)$ cosmology. The equivalent scalar field description is also reported in detail. The dynamical consequences, adopting the modified Friedmann equations, are reported in Sect. III. There, we also baptize our dynamical variables and set the system of differential equations, used throughout our work. The consequences on our observable universe are summarized in Sect. IV. There, we report the cases of spatially-flat and non-flat universe and the overall analysis. Finally, the scalar field potentials and their impact under the form of power-law contribution is explored in Sect. V, whereas our final outlooks and perspectives are summarized in Sect. VI.
II $f\left(G\right)$-Cosmology
At large scales, the maximally symmetric spacetime describing the universe is the four-dimensional isotropic and homogeneous Friedmann-Lemaitre-Robertson-Walker geometry,
$ds^{2}=-N^{2}\left(t\right)dt^{2}+a^{2}\left(t\right)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right)\right],$ (1)
with $a\left(t\right)$ representing the cosmic scale factor, $N\left(t\right)$ the lapse function and $k$ denotes the spatial curvature of the associated three-dimensional hypersurface.
The latter indicates the universe topology, namely for $k=0$, the universe is spatially flat, while for $k=-1$, the hypersurface has negative curvature, turning out to be hyperbolic and, finally, $k=1$ leads to an hypersurface represented by a closed sphere.
For the gravitational theory, we assume the modified four-dimensional Gauss-Bonnet theory with the action, $S[R,G]=\int d^{4}x\sqrt{-g}\,\mathcal{L}[R,G]$, given by Li:2007jm
$S[R,G]=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+f\left(G\right)\right),$ (2)
where $f\left(G\right)$ is an arbitrary function of the Gauss-Bonnet scalar, $G$, and $R$ is the Ricci scalar, providing standard Einstein’s gravity.
As stated in the introduction, in a four-dimensional manifold, the Gauss-Bonnet scalar is a topological invariant, i.e., $G$ represents a boundary term, while considering any linear $f\left(G\right)$ function, de facto lets Eq. (2) turn into the equivalent Hilbert-Einstein action.
Accordingly, we hereafter focus on non-linear $f\left(G\right)$ contributions, with the aim of exploring its effects over the cosmic expansion history.
II.1 Scalar field equivalent description
Remarkably, introducing a Lagrange multiplier would easily provide an equivalent description in terms of scalar fields. Accordingly, we can show that the geometrodynamical degrees of freedom provided by the nonlinear function $f\left(G\right)$ can be mimicked with a scalar field [^1] A scalar field is commonly associated with interactions and particles. It is always possible, however, to attribute dynamical properties of a given geometrical model to it, with the great advantage of handling a representation whose stability is easier to compute..
Specifically, the gravitational model in Eq. (2) is equivalent to the Einstein-Gauss-Bonnet-Scalar field theory, where the scalar field is coupled to $G$.
Motivated by this recipe, we employ the Lagrange multiplier, $\lambda$, and write the Eq. (2) as
$S_{f\left(G\right)}=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+f\left(G\right)-\lambda\left(G-\mathcal{G}\right)\right),$ (3)
where, solving the equations of motion, leads to the constraint $G=\mathcal{G}$, that will be essential for our computation later on.
By varying the Lagrangian with respect to the Gauss-Bonnet term gives
$\lambda=f_{,G}\left(G\right),$ (4)
that permits to reformulate the Lagrangian as
$\mathcal{L}=\frac{R}{2}+\left(f\left(G\right)-Gf_{,G}\left(G\right)\right)+f_{,G}\left(G\right)\mathcal{G}.$ (5)
Conveniently, ensuring that $f(G)$ and $f_{,G}(G)$ are distinct variables, we can introduce the scalar field representation by
$\phi=f_{,G}\left(G\right),$ (6a)
$V\left(\phi\right)=\left(f\left(G\right)-Gf_{,G}\left(G\right)\right),$ (6b)
ending up with the equivalent Einstein-Gauss-Bonnet scalar field picture
$S_{f\left(G\right)}=\int d^{4}x\sqrt{-g}\left(\frac{R}{2}+\phi G+V\left(\phi\right)\right).$ (7)
This is the Action Integral for the Einstein-Gauss-Bonnet Scalar field gravitational model Fomin:2018typ ; Odintsov:2020xji ; Paliathanasis:2024gwp ; Millano:2024vju ; TerenteDiaz:2023iqk ; Nojiri:2023jtf ; Hussain:2024yee without a kinetic term.
Afterwards, from the line element in Eq. (3), the Ricci, $R$, and the Gauss-Bonnet, $G$, scalars are
$R=6\left(\frac{1}{N}\dot{H}+2H^{2}+\frac{k}{a^{2}}\right),$ (8a)
$G=24\left(\frac{1}{N}\dot{H}+H^{2}\right)\left(H^{2}+\frac{k}{a^{2}}\right),$ (8b)
where we consider the Hubble function to be, $H\equiv\frac{1}{N}\frac{\dot{a}}{a}$.
Hence, plugging the Ricci and Gauss-Bonnet scalars into Eq. (7) and, thus, integrating by parts, we obtain the following point-like Lagrangian,
$\mathcal{L}\left(N,a,\dot{a},\phi,\dot{\phi}\right)=\frac{3}{N}a\dot{a}^{2}-\frac{8}{N^{3}}\dot{a}^{3}\dot{\phi}+3ka\left(N-\frac{8}{N}\frac{\dot{a}}{a}\dot{\phi}\right)+Na^{3}V\left(\phi\right),$ (9)
from which it appears quite natural to compute the cosmological field equations with respect to the dynamical variables $\left\{N,a,\phi\right\}$. Specifically, the variation with respect to the lapse function, $N$, gives the constraint equation.
$3H^{2}-24H^{3}\frac{\dot{\phi}}{N}-\frac{3k}{a^{2}}\left(1+8H\frac{\dot{\phi}}{N}\right)-V\left(\phi\right)=0,$ (10)
while varying with respect to the scale factor, $a$, and to the scalar field, $\phi$, provides the following second-order (for $a(t)$ and $\phi$) differential equations,
$\frac{1}{N}\dot{H}+H^{2}+\frac{V_{,\phi}}{24\left(H^{2}+\frac{k}{a^{2}}\right)}=0,$ (11a)
$\frac{12\ddot{\phi}}{N^{2}}\left(\frac{k}{a^{2}}+H^{2}\right)^{2}-3\left(\frac{k}{a^{2}}\right)^{2}+\left(3H^{4}+\frac{V_{,\phi}}{8}\right)-12\left(\frac{k}{a^{2}}+H^{2}\right)^{2}\left(\frac{3}{N}H+\frac{\dot{N}}{N^{3}}\right)-\frac{1}{8}\frac{\dot{\phi}}{N}HV_{,\phi}=0.$ (11b)
Bearing these results in mind, we can accordingly compute the cosmological dynamics.
III Cosmological dynamics
We here work out a detailed phase-space analysis based on Eqs. (10), (11a) and (11b).
Enabling the Hubble rate to space over positive and negative regions, we can introduce the dimensionless variables Paliathanasis:2024gwp ; Millano:2024vju
$x=8\frac{\dot{\phi}}{1+H^{2}},$ (12a)
$y=\frac{V\left(\phi\right)}{1+H^{2}},$ (12b)
$\eta=\frac{H}{\sqrt{\left(1+H^{2}\right)}},$ (12c)
$\omega_{k}=\frac{k}{a^{2}\left(1+H^{2}\right)},$ (12d)
$\lambda=\frac{V_{,\phi}}{V},$ (12e)
$d\tau=(1+H^{2})dt,$ (12f)
which differ from the $H$-normalization approach Copeland:2006wr since the assumption made through these parameters allows us to pass through $H=0$, i.e., to let $H$ change sign, as stated above.
Selecting $\tau$ as the new independent variable, the field equations turn into
$\left(\eta^{4}+\left(1-\eta^{2}\right)\omega_{k}\right)^{2}\frac{dx}{d\tau}$ $=-\frac{y\eta^{2}}{24}\left(1-\eta^{2}\right)\left(\eta^{2}\left(2\lambda-\eta\left(2\left(\lambda-12\right)\eta+\lambda x\left(2+\eta^{2}\right)\right)\right)\right)$
$-\frac{y\eta^{2}}{24}\left(1-\eta^{2}\right)^{2}\left(\lambda x\eta-24\right)\omega_{k}^{2}+\omega_{k}^{4}\left(1-\eta^{2}\right)^{2}\left(\eta^{2}\left(\eta x-1\right)-1\right)$
$+\left(\eta^{7}\left(\eta\left(1+\eta^{2}\left(\eta x-1\right)\right)-2x\left(1-\eta^{2}\right)\omega_{k}^{2}\right)\right),$ (13)
$\frac{dy}{d\tau}=-\frac{1}{48}y\sqrt{1-\eta^{2}}\left(3\lambda x+2\eta^{3}\left(24+\frac{\lambda y\left(1-\eta^{2}\right)}{\eta^{4}-\left(1-\eta^{2}\right)\omega_{k}}\right)\right),$ (14)
$\frac{d\eta}{d\tau}=\frac{y}{24\eta}\left(3\lambda x\eta+48\eta^{4}+2\lambda y\left(1-\eta^{2}\right)\right),$ (15)
$\frac{d\omega_{k}}{d\tau}=\frac{\left(1-\eta_{k}\right)^{2}}{12\eta}\omega_{k}^{2}\left(24+\frac{\lambda y\left(1-\eta^{2}\right)}{\eta^{4}-\left(1-\eta^{2}\right)\omega_{k}}\right),$ (16)
and
$\frac{d\lambda}{d\tau}=\frac{\lambda}{8}x\left(\Gamma\left(\lambda\right)-1\right),$ (17)
where $\Gamma\left(\lambda\right)=\frac{V_{,\phi\phi}V}{\left(V_{,\phi}\right)^{2}}$.
Also, from Eq. (10) the algebraic constraint becomes
$\frac{1}{3}y+\omega_{k}-\eta^{2}+\frac{\eta}{1-\eta^{2}}x\left(\eta^{2}+\omega_{k}\right)=0.$ (18)
By construction, $\eta$ is limited as $\eta^{2}\leq 1$, although the other variables are not compactified. Indeed, for $\eta^{2}=1$, the solution describes a scaling universe, while for $\eta^{2}<1$, the solution describes a de Sitter solution. When $\eta>0$ the dynamical system describes an expanding universe $\left(H>0\right)$, while for $\eta<0$, the dynamical system describes a collapsing universe $\left(H<0\right)$.
In terms of the dimensionless variables the deceleration parameter $q=-1-\frac{\dot{H}}{H^{2}}$ 2016IJGMM..1330002D ; 2012PhRvD..86l3516A ; 2014PhRvD..90d3531A ; 2017PDU….17…25A ; 2024A&A…690A..40L ; 2020A&A…641A.174L is expressed as
$q\left(x,y,\eta,\omega_{k}\right)=\frac{\lambda}{24}\frac{1-\eta^{2}}{\eta^{2}}\frac{y}{\eta^{2}+\omega_{k}}.$ (19)
Immediately we notice that,
    The field equations form a five-dimensional dynamical system of first-order differential equations. By using the constraint provided in Eq. (18), the dimension of the dynamical system is reduced by one.
    Moreover, for the exponential potential $V\left(\phi\right)=V_{0}e^{\lambda_{0}\phi}$, it follows that $\Gamma\left(\lambda\right)-1=0$, thus, parameter $\lambda$ is always a constant, i.e. $\lambda=\lambda_{0}.$
    Hence, in order to reduce further the dimension of the dynamical system we consider the exponential potential.
    Indeed, for the exponential potential $V\left(\phi\right)$, the corresponding $f\left(G\right)$ function is derived
    $f\left(G\right)=\frac{G}{\lambda_{0}}\left[\ln\left(-\frac{G}{\lambda_{0}V_{0}}\right)-1\right].$ (20)
IV Consequences on the observable Universe
We can now wonder whether our scenario can be matched with the observable universe. This would be important in order to check the goodness of our approach and to infer limits about the initial Lagrangian, proposed in Eq. (3). To do so, we focus below on spatially-flat and non-flat universe and find the possible attractors and stability at infinity.
IV.1 Spatially-flat universe
Consider now a spatially flat FLRW universe, $k=0$, rewriting $y$ from Eq. (18), the two-dimensional system yields, $\left\{x,\eta\right\}$, that is,
$\frac{dx}{d\tau}=1-\eta^{2}\left(1-x\eta\right)-\frac{\left(\eta\left(x+\eta\right)-1\right)\left(\lambda x\eta\left(2+\eta^{2}\right)-2\lambda+2\eta^{2}\left(\lambda-12\right)\right)}{8\eta^{2}},$ (21)
$\frac{d\eta}{d\tau}=-\frac{1}{24}\eta^{2}\left(1-\eta^{2}\right)\left(24+\frac{\lambda\eta^{2}\left(1-\eta^{2}\right)\left(1-x\eta+\eta^{2}\right)}{1-\eta^{2}}\right).$ (22)
The stationary/equilibrium points $A=\left(x\left(A\right),\eta\left(A\right)\right)$ of the latter two-dimensional system are
$A_{1}$ $=\left(0,1\right),~{}A_{2}=\left(0,-1\right),~{}$
$A_{3}$ $=\left(\frac{32}{3}\frac{1}{\lambda},1\right)~{},~{}A_{4}=\left(-\frac{32}{3}\frac{1}{\lambda},-1\right),$
$A_{5}$ $=\left(0,\sqrt{\frac{\lambda}{\lambda-8}}\right),~{}A_{6}=\left(0,-\sqrt{\frac{\lambda}{\lambda-8}}\right).$
The equilibrium points with odd numbers describe expanding universes, while the equilibrium points with even numbers describe collapse. Points $A_{5}$, $A_{6}$ describe de Sitter solutions, while the remaining points describe scaling exact solutions.
[…]
Equilibrium points $A_{3}$ and $A_{4}$ describe scaling solutions which can describe cosmic acceleration, that is, $q\left(A_{3}\right)=-\frac{4}{3}$ and $q\left(A_{4}\right)=-\frac{4}{3}$. At these points parameter $y$ reaches infinity. The eigenvalues for point $A_{3}$ are $\left\{-4,-\frac{2}{3}\right\}$, while for point $A_{4}$ are $\left\{4,\frac{2}{3}\right\}$. Therefore, point $A_{3}$ is always an attractor and point $A_{4}$ is always a source point.
Furthermore, equilibrium points $A_{5}$ and $A_{6}$ describe de Sitter solutions with $q\left(A_{5}\right)=-1$, and $q\left(A_{6}\right)=-1$. The points are real for $\lambda<0.$ At these points there is a nonzero contribution of the scalar field potential, that is, $y=\frac{3\lambda}{\lambda-8}$. The eigenvalues of the linearized systems around these points are $\left\{-\frac{3+\sqrt{17}}{2}\sqrt{\frac{\lambda}{\lambda-8}},-\frac{3-\sqrt{17}}{2}\sqrt{\frac{\lambda}{\lambda-8}}\right\}$, from where we conclude that the equilibrium points are always saddle points.
IV.1.1 Analysis at infinity
Above, we computed the equilibrium points for the dynamical system, reported in Eqs. (21) and (22), in the finite regime.
Nevertheless, clarifying the asymptotic regime appears quite interesting, as the variable $x$ can acquire values at infinity.
Accordingly, we thus introduce the new compactified variable, $X$, so that $x=\frac{X}{\sqrt{1-X^{2}}}$ and study the existence of equilibrium points at infinity, where $X^{2}=1$.
Indeed, there are four points, $\hat{A}=\left(X\left(\hat{A}\right),\eta\left(\hat{A}\right)\right),$ with coordinates
$\hat{A}_{1}^{\pm}=\left(\pm 1,1\right)~{},~{}\hat{A}_{2}^{\pm}=\left(\pm 1,-1\right).$
At these points we calculate the deceleration parameter $q\left(\hat{A}_{1}^{\pm}\right)=-sign\left(\lambda X\left(\hat{A}_{1}^{\pm}\right)\right)\infty$ and $q\left(\hat{A}_{2}^{\pm}\right)=sign\left(\lambda X\left(\hat{A}_{2}^{\pm}\right)\right)\infty$. Points $\hat{A}_{1}^{\pm}$ describe an expanding universes, but for $sign\left(\lambda X\left(\hat{A}_{1}^{\pm}\right)\right)<0,$ we derive that $q\left(\hat{A}_{1}^{\pm}\right)>0$, that means these stationary points correspond to asymptotic solutions, asymptotically reaching the Minkowski space.
The same holds for the points exhibiting $sign\left(\lambda X\left(\hat{A}_{2}^{\pm}\right)\right)<0$.
Consequently, the existence of these stationary points is essential in order the cosmological evolution to cross from expansion to collapse and vice versa. Recall that $\eta=0$, gives a singularity at the dynamical system, so the transition is happening through the points at infinity.
This prescription is explicitly reported in Fig. 1. Last but not least, the stationary points at infinity always correspond to unstable solutions.
From the phase-space portraits displayed in Fig. 1, it appears clear that the most of underlying initial conditions leads to a collapsed universe, whereas only a small fraction of the entire phase-space, less than one quarter, corresponds to trajectories with initial conditions leading to a future attractor and, then, describing cosmic acceleration.
Figure 1: Phase-space portrait for the dynamical system (21), (22) in the compactified variables $X$ and $\eta$. Equilibrium points at the infinity are marked with red, the source points $A_{1}$ and $A_{4}$ are marked with orange, the saddle points $A_{5}$ and $A_{6}$ with green, and the attractors $A_{2}$ and $A_{5}$ are marked with blue. We observe that the main of the initial conditions lead to a collapsed universe, while there exist only a small area, less than then one quarter, of the phase-space, where the trajectories lead to a future attractor which describes cosmic acceleration.
A summary of the above outcomes is summarized in Table 1.
Table 1: Equilibrium Points for the spatially flat FLRW universe and exponential potential.
Point $\left(\mathbf{x,\eta}\right)$ Existence $q$ Stable?
$A_{1}$ $\left(0,1\right)$ Always 00 False
$A_{2}$ $\left(0,-1\right)$ Always 00 Attractor
$A_{3}$ $\left(\frac{32}{3}\frac{1}{\lambda},1\right)$ $\lambda\neq 0$ $-\frac{4}{3}$ Attractor
$A_{4}$ $\left(-\frac{32}{3}\frac{1}{\lambda},-1\right)$ $\lambda\neq 0$ $-\frac{4}{3}$ False
$A_{5}$ $\left(0,\sqrt{\frac{\lambda}{\lambda-8}}\right)$ $\lambda<0$ $-1$ False
$A_{6}$ $\left(0,-\sqrt{\frac{\lambda}{\lambda-8}}\right)$ $\lambda<0$ $-1$ False
$\hat{A}_{1}^{\pm}$ (±$\left(\pm\infty,1\right)$ Always $\mp sign\left(\lambda\right)\infty$ False
$\hat{A}_{2}^{\pm}$ (±$\left(\pm\infty,-1\right)$ Always ±$\pm sign\left(\lambda\right)\infty$ False
IV.2 Non-zero spatial curvature
Let us now assume now that the spatial curvature does not vanish. Accordingly, the variable $\omega_{k}$ furnishes a non-trivial dynamics.
We determine the equilibrium points $B=\left(x\left(B\right),\eta\left(B\right),\omega_{k}\left(B\right)\right)$ for the dynamical system in Eqs. (13), (15) and (16), reading
$B_{1}$ $=\left(0,1,\omega_{k}\right),~{}B_{2}=\left(0,-1,\omega_{k}\right),$
$B_{3}$ $=\left(\frac{8}{3\lambda}\frac{4+3\omega_{k}}{1+3\omega_{k}},1,\omega_{k}\right),$
$B_{4}$ $=\left(-\frac{8}{3\lambda}\frac{4+3\omega_{k}}{1+3\omega_{k}},-1,\omega_{k}\right),$
$B_{5}$ $=\left(0,\sqrt{\frac{\lambda}{\lambda-8}},0\right),~{}B_{6}=\left(0,-\sqrt{\frac{\lambda}{\lambda-8}},0\right),$
where at points $B_{1},~{}B_{2},~{}B_{3}$ and $B_{4}$$\omega_{k}$ is undefined since these do not represent just points, but surfaces defined on the parametric space with varying $\omega_{k}$. Hence, the solution trajectories lie on these surfaces and move along $\omega_{k}$.
The equilibrium points $B_{1},~{}B_{2}$ describe cosmic expansion and collapse respectively. The corresponding deceleration parameters are $q\left(B_{1}\right)=0$ and $q\left(B_{2}\right)=0$. The eigenvalues, computed near $B_{1}$, are $\left\{2,4+3\omega_{k},0\right\}$, while near $B_{2}$, they are $\left\{-2,-4-3\omega_{k},0\right\}$. The equilibrium point $B_{1}$ is always unstable, differently from $B_{2}$, for which there can be a stable submanifold.
To check this, we expand up to the second-order approximation, seeking the $\omega_{k}$ perturbations, and find the exact solution for the evolution of the perturbations $\delta x\simeq e^{-2t}+\alpha_{0}e^{-\left(4+3\omega_{k}\right)t}$$\delta\eta\simeq e^{-2t}$ and $\delta\omega_{k}\simeq e^{-4t}+\alpha_{1}e^{-3\left(2+\omega_{k}\right)t}$ from which we conclude that $B_{2}$ turns out to be stable.
Furthermore, the equilibrium points $B_{3}$ and $B_{4}$ are defined for $\omega_{k}\neq-\frac{1}{3}$ and appear the generalizations of $A_{3}$ and $A_{4}$, respectively. The deceleration parameters are $q\left(B_{3}\right)=-1-\frac{1}{3\left(1+\omega_{k}\right)}$ and $q\left(B_{4}\right)=-1-\frac{1}{3\left(1+\omega_{k}\right)}$. Thus, here the cosmic acceleration is recovered either for $\omega_{k}<-\frac{4}{3}$ or $\omega_{k}>-1$.
The eigenvalues of the linearized system around the point $B_{3}$ are $\left\{-4-3\omega_{k},-\frac{2}{3}\left(1+3\omega_{k}\right),0\right\},$ while at the point $B_{4}$ we find $\left\{4+3\omega_{k},\frac{2}{3}\left(1+3\omega_{k}\right),0\right\}$.
In analogy to what found before, we study the evolution of the perturbations up to the second-order and it follows that $B_{3}$ appears as an attractor for $\omega_{k}>-\frac{1}{3}$, while $B_{4}$ is an attractor with $\omega_{k}<-\frac{4}{3}$.
Finally, the stationary points $B_{5},$ $B_{6}$ describe a de Sitter solution of the spatially flat universe of the points $A_{5}$ and $A_{6}$.
Last but not least we found that $B_{5}$, $B_{6}$ share the two eigenvalues with points $A_{5}$ and $A_{6}$, from which we infer that they are always saddle points.
IV.2.1 Analysis at infinity
We employ the same compactified variable $X$ as before, and we find that the stationary points at the infinity are
$\hat{B}_{1}^{\pm}=\left(\pm 1,1,\omega_{k}\right)~{},~{}\hat{B}_{2}^{\pm}=\left(\pm 1,-1,\omega_{k}\right).$
They provide the same physical properties as in the flat case, while remarkably the stationary points always describe unstable solutions.
If we consider that $\left|\omega_{k}\right|$ is small, then the future attractors are the points $B_{2}$ and $B_{3}$.
These points may exhibit nonzero curvature. However, since the phase-space trajectories pass through a de Sitter expansion, in which $\eta^{2}<1$, then the parameter $\left|\omega_{k}\right|$ tends to decrease, when the attractor is point $B_{3}$. This may be reinterpreted as a solution of the flatness problem. In order to demonstrate this, consider that we start from initial conditions of a nonzero curvature FLRW space. Then the trajectories will move along the surfaces we described before. Parameter $10<\eta<$ 1, gives necessary a de Sitter expansion, since for $\eta_{0}$, $H=\frac{\eta_{0}}{\sqrt{1-\eta_{0}^{2}}}$. Then $\left|\omega_{k}\right|=\left|\frac{k}{a^{2}\left(1+H^{2}\right)}\right|$, $\downarrow$, because $a\uparrow$, then $\omega_{k}\rightarrow 0$, asymptotically. .
Otherwise in the collapse, i.e., as the point $B_{2}$ is the attractor parameter, then $\left|\omega_{k}\right|$ tends to increase. However, information toward the original sign of curvature of the dynamical system remains unaltered at the attractors.
The phase-space behavior is thus analogous to that displayed in Fig. 1, where it turns out that the most initial conditions result into a gravitational collapse.
In term of dynamics the curvature component can be seen equivalently in the same form of a spatially flat geometry induced with a indeal gas with constant equation of state $p_{m}=-\frac{1}{3}\rho_{m.}$. By using this analogue, we can compare our results with the matter dominated investigated in Goheer:2009qh . Nevertheless this is an asymptotic solution, near a specific value of the Gauss-Bonnet scalar because $x=0$, that is, the function (20 ) is described by a polynomial similar to that determined in Goheer:2009qh .
V Fixing the scalar field potential
The use of exponential scalar field, as above reported, is quite usual in the literature Paliathanasis:2024gwp ; Millano:2024vju ; Copeland:2006wr ; Carloni:2023egi . The advantage of using such a potential is to reduce the complexity of variables under exam. However, more appropriate scalar field potentials may be even associated with dark energy scenarios, see e.g. 2024arXiv241010935C ; Lazkoz:2007mx ; Paliathanasis:2024jxo ; Papagiannopoulos:2016dqw and, then, turn out to be relevant to check the validity of our background theory. Below, we employ a class of power law potentials, with unfixed exponents, to explore the impact in our stability treatment.
V.1 The role of power-law potential
All our treatment made use of the exponential potential, for which Eq. (17) turns out to be trivially satisfied, yielding a constant $\lambda$.
For completeness, it appears essential to explore alternative potentials, in order to evaluate the impact to dynamical variables, within the usual context of a spatially flat FLRW geometry.
The simplest approach is based on ensuring the validity of a power-law potential,
$V\left(\phi\right)=V_{0}\phi^{n},$ (23)
from where it follows $\lambda=\frac{n}{\phi}$ and $\Gamma\left(\lambda\right)=1-\frac{1}{n}$, obtained by definition, since $\Gamma\left(\phi\right)=\frac{n\left(n-1\right)\phi^{n-2}\phi^{n}}{n^{2}\left(\phi^{n-1}\right)^{2}}=\frac{n-1}{n}$.
In view of Eq. (23), Eq. (17) furnishes
$\frac{d\lambda}{d\tau}=\frac{1}{8n}\lambda x.$ (24)
Remarkably, for this potential we calculate the power-law $f\left(G\right)$ model, that approximately reads $f\left(G\right)\simeq G^{\frac{n}{n-1}}$.
The equilibrium points $C=\left(x\left(C\right),\eta\left(C\right),\lambda\left(C\right)\right)$ of the dynamical system, namely Eqs. (21), (22) and (24), are
$C_{1}$ $=\left(A_{1},\lambda\right),~{}C_{2}=\left(A_{2},\lambda\right),~{}$
$C_{3}$ $=\left(A_{5},\lambda\right),~{}C_{4}=\left(C_{6},\lambda\right),$
where the parameter $\lambda$ above, for all these points, appears arbitrary. Phrasing it differently, the points actually describe a family of solutions for different values of $\lambda$. In the stationary points $\lambda$ is varying without affective the solution trajectories, in a similar way as before for the $\omega_{k}$. When $\phi\rightarrow\infty$, for $n>0$, or $\phi\rightarrow 0$, for $n<0$, the potential term dominates, and there is not any contribution from the potential term, i.e. $x=0$, that exactly is happening here. The value of the parameter $\lambda$ is defined by the initial conditions.
We remark that all the equilibrium points have $x=0$.
[…]
Point $C_{1}$ is found to always be a source, while point $C_{2}$ is an attractor.
For the equilibrium point $C_{4}$ we calculate the eigenvalues,
$\left\{0,\frac{3\sqrt{n\lambda}-\sqrt{17n\lambda+8}}{2\sqrt{n\left(\lambda-8\right)}},\frac{3\sqrt{n\lambda}+\sqrt{17n\lambda+8}}{2\sqrt{n\left(\lambda-8\right)}}\right\},$ (25)
from where conclude that the point describes an unstable solution. Specifically it is a saddle point for $n>-\frac{1}{\lambda}$, provided that the existence condition requires $\lambda<0$.
Finally, point $C_{3}$ describes a de Sitter expanding universe and the corresponding eigenvalues are
$\left\{0,\frac{-3\sqrt{n\lambda}-\sqrt{17n\lambda+8}}{2\sqrt{n\left(\lambda-8\right)}},\frac{-3\sqrt{n\lambda}+\sqrt{17n\lambda+8}}{2\sqrt{n\left(\lambda-8\right)}}\right\}.$ (26)
[…]
V.2 Asymptotic analysis at Infinity
Equilibrium points at infinity exist only as $\lambda=0$ and their physical properties are those described by $\hat{A}^{\pm}$.
We end up that none of the equilibrium point at infinity is an attractor.
In Fig.2, we present three-dimensional phase-portraits for two values of parameter $n$ and remark that, for $n>0,$ there exists a region where the trajectories reach on the stable surface of de Sitter expanding solutions, described by the family of points $C_{3}$.
Figure 2: Phase-space portrait for the dynamical system (21), (22) and (24) in the variables $x,\eta~{}$and $\lambda$ where the family of points $C_{3}$ is marked. Left Fig. is for $n=\frac{1}{5}$ and right Fig. for  $n=-\frac{1}{5}$ It is easy to see that for $n=\frac{1}{5}$, the trajectories lie on the stable surface $C_{3}$, and the de Sitter solution is an attractor.
V.3 Discussion for a general potential
Consider now the generic nonlinear potential $V\left(\phi\right)$, that is, a generic nonlinear function $f\left(G\right)$. Then the equilibrium points solve the algebraic equation (17), that is,
$\frac{\lambda}{8}x\left(\Gamma\left(\lambda\right)-1\right)=0$
that is, $x=0$ or $\lambda=\lambda_{0}$ such that $\lambda_{0}\left(\Gamma\left(\lambda_{0}\right)-1\right)=0$, where $\Gamma\left(\lambda\right)$ is constructed by the potential.
When $x=0$, it follows from before that $\lambda$ does not constrained by the rest equations of the dynamical system, that is, the admitted equilibrium points are those described by the power-law potential.
On the other hand, for $\lambda=\lambda_{0}$, then from the rest of the equations variables $x,\eta$ are constrained and provide families of equilibrium points as given by the exponential potential, for each value of $\lambda_{0}$.
Therefore, the definition of an arbitrary potential function will not lead to the derivation of new families of solutions. Hence, the study of the exponential and the power-law potentials is sufficient to construct all the possible asymptotic solutions.
However, what is important to mention is that the stability properties of the equilibrium points will not remain the same. Because the function $\Gamma\left(\lambda\right)$ and, its derivative play an important role in the linearization of the dynamical system around the equilibrium points.
VI Final remarks
In this work, we explored a field-equivalent representation of $f(G)$-theories of gravity, with the aim of investigating the phase-space dynamics and the corresponding stability. To do so, we specifically considered a FLRW universe with flat topology and introduced a Lagrange multiplier within the modified Lagrangian density, de facto reformulating the action of $f(G)$-gravity into an equivalent form of Einstein-Gauss-Bonnet scalar field cosmology, that does not exhibit an explicit scalar field kinetic term.
The scalar field potential is directly related to the function $f(G)$ and its derivatives. Further, within this scalar field formulation, we computed the cosmological field equations and introduced a set of dimensionless variables, quite different from those used in standard approach, i.e., carrying out a global dynamical analysis by enabling the Hubble rate to change its sign.
In view of the fact that the field equations constituted a complicated constrained dynamical system, by utilizing the constraint equation, we reduced the system’s dimensionality by one, reducing to four. In addition, we showed that, using an exponential potential, the dynamical system is further reduced by another dimensionality. Afterwards, we identified the equilibrium points and found two attractors, both independent from the spatial curvature.
Within our analysis, one equilibrium point described an expanding, accelerating universe, not behaving as a de Sitter solution, while another one corresponded to a collapsing spacetime.
More precisely, even though a de Sitter solution exists, it manifests as a saddle point. Nevertheless, the phase-space analysis indicated that most initial conditions lead to a collapsed universe, pointing out that our scenario is likely unphysical.
Indeed, in the power-law $f(G)$-theory, the scaling solution appears absent, and the de Sitter solution is a future attractor for positive values of the power parameter and for specific set of initial conditions. That means that while the $f(G)$-theory can provide an explanation to the cosmic acceleration, the majority of the trajectories of the solutions lead to attractors which can not describe the present cosmic acceleration. Nevertheless, unstable equlibrium points which describe cosmic acceleration can be related to the early-time inflation.
Future works will try to extend our treatment to more physical contexts in which the topological terms may imply more stable regions, whose physical sense can be more defined. In particular, by adding further external fields, such as sources to the energy momentum-tensor or trying to include additional constraints. The impact of perturbations will be also investigated, in order to fix more carefully the free parameters of the theory, as well as possible additional potentials, based on precise physical meanings, related to dark energy sources 2024arXiv241010935C or inflationary domains.
Acknowledgements.
GL and AP are grateful for the support of Vicerrectoría de Investigación y Desarrollo Tecnológico (Vridt) at Universidad Católica del Norte through Núcleo de Investigación Geometría Diferencial y Aplicaciones, Resolución Vridt No - 096/2022 and Resolución Vridt No - 098/2022. GL and AP were economically supported by the Proyecto Fondecyt Regular 2024, Folio 1240514, Etapa 2024. AP expresses his gratitude to the University of Camerino and to OL for invitation and hospitality provided during the time in which this work has been written. OL acknowledges financial support from the Fondazione ICSC, Spoke 3 Astrophysics and Cosmos Observations. National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) Project ID CN__\__00000013 ”Italian Research Center on High-Performance Computing, Big Data and Quantum Computing” funded by MUR Missione 4 Componente 2 Investimento 1.4: Potenziamento strutture di ricerca e creazione di ”campioni nazionali di R&\&&S (M4C2-19 )” - Next Generation EU (NGEU) GRAB-IT Project, PNRR Cascade Funding Call, Spoke 3, INAF Italian National Institute for Astrophysics, Project code CN00000013, Project Code (CUP): C53C22000350006, cost center STI442016.